# Direct Preference Optimization Needs

### v3

@@ Abstract
Aligning large language models (LLMs) with human values and intentions is crucial for their utility, honesty, and safety. Reinforcement learning from human feedback (RLHF) is a popular approach to achieve this alignment, but it faces challenges in computational efficiency and training stability. Recent methods like Direct Preference Optimization (DPO) and Simple Preference Optimization (SimPO) have proposed offline alternatives to RLHF, simplifying the process by reparameterizing the reward function. However, DPO depends on a potentially suboptimal reference model, and SimPO’s assumption of a fixed target reward margin may lead to suboptimal decisions in diverse data settings. In this work, we propose $\alpha$-DPO, an adaptive preference optimization algorithm designed to address these limitations by introducing a dynamic reward margin. Specifically, $\alpha$-DPO employs an adaptive preference distribution, balancing the policy model and the reference model to achieve personalized reward margins. We provide theoretical guarantees for $\alpha$-DPO, demonstrating its effectiveness as a surrogate optimization objective and its ability to balance alignment and diversity through KL divergence control. Empirical evaluations on AlpacaEval 2 and Arena-Hard show that $\alpha$-DPO consistently outperforms DPO and SimPO across various model settings, establishing it as a robust approach for fine-tuning LLMs. Our method achieves significant improvements in win rates, highlighting its potential as a powerful tool for LLM alignment. The code is available at https://github.com/junkangwu/alpha-DPO.
@@ 1 Preliminaries
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right].$ (4)
@@
${\mathcal{L}}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$ (5)
@@ 2 Method
In this section, we establish a unified framework that connects DPO and SimPO (Section 2.1), highlighting the critical role of the reference model in preference optimization. We then introduce $\alpha$-DPO (Section 2.2), a novel preference optimization algorithm that synergizes the strengths of both DPO and SimPO.
@@ 2.1 A Common Framework for DPO and SimPO
$\mathcal{L}(\pi_{\theta};U)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\pi_{\theta}(y_{w}|x)-\log\pi_{\theta}(y_{l}|x)\right)-\gamma\right)\right],$ (6)
@@
$\mathcal{L}_{\text{LN}}(\pi_{\theta};U)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right].$ (7)
@@
Theorem 2.1 establishes a unified framework for DPO and SimPO by showing that replacing the reference model $\pi_{\text{ref}}$ in DPO with a uniform distribution $U$ reduces the DPO loss to the SimPO loss, up to a constant term $\gamma$. This reveals that SimPO is essentially DPO with a uniform reference model. Consequently, the term $\beta\left(\log\pi_{\text{ref}}(y_{w}|x)-\log\pi_{\text{ref}}(y_{l}|x)\right)$ collapses to a constant, emphasizing the pivotal role of the implicit reference model in preference optimization.
Figure 1: $\log\pi_{\mathrm{ref}}(y_{w}|x)-\log\pi_{\mathrm{ref}}(y_{l}|x)$ along the traiing steps.
@@
Limitations of SimPO: While SimPO simplifies the loss function by using a constant offset $\gamma$, it overlooks the variability inherent in different data instances, as $\gamma$ remains the same across all training samples. This rigidity could lead to suboptimal performance, especially in the presence of noise or inconsistencies in the data. Moreover, completely discarding the original reference model $\pi_{\text{ref}}$ may fail to capture important distinctions between response pairs that could be leveraged to improve learning.
2.2 Proposed Method: $\alpha$-DPO
@@
Deriving the $\alpha$-DPO Objective. Starting from the standard Reinforcement Learning (RL) objective for preference optimization, we redefine the reference model $\pi_{\text{ref}}$ as an implicit reference model $\hat{\pi}_{\text{ref}}$, formulated as:
$\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\alpha},$ (8)
where $\alpha$ is a hyperparameter controlling the influence of the policy model on the reference model, and $U(y|x)$ is a uniform distribution serving as a constant baseline. When $\alpha=0$, $\hat{\pi}_{\text{ref}}$ reduces to the uniform distribution as in SimPO; when $\alpha=1$, it incorporates the ratio between the policy and reference models as in DPO.
Substituting $\hat{\pi}_{\text{ref}}$ into the original DPO loss function, we obtain the $\alpha$-DPO objective:
$\mathcal{L}_{\text{$\alpha$-DPO}}(\pi_{\theta},\pi_{\text{ref}})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}-\beta\log\frac{\hat{\pi}_{\text{ref}}(y_{w}|x)}{\hat{\pi}_{\text{ref}}(y_{l}|x)}\right)\right]$ (9)
$=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right)-\left[\gamma+\alpha M(x,y_{w},y_{l})\right]\right)\right],$
@@
Length-normalized Reward Formulation: Inspired by the technique used in SimPO, we incorporate length normalization into our method. This adjustment ensures that rewards are scaled appropriately with respect to the length of the sequences, thereby stabilizing the training process. As demonstrated in our experiments (cf. Appendix B.2), we also confirmed that even without length normalization, our method remains effective and continues to show performance improvements.
@@
Final Objective: Incorporating the above considerations, the final $\alpha$-DPO loss function becomes:
$\mathcal{L}_{\text{$\alpha$-DPO}}(\pi_{\theta},\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(u(x,y_{w},y_{l})-\operatorname{sg}\left[\gamma+\alpha M^{\ast}(x,y_{w},y_{l})\right]\right)\right],$ (12)
@@
where $u(x,y_{w},y_{l})=\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log{\pi_{\theta}(y_{l}|x)}$. This formulation ensures balanced influence between the policy and reference models, aligning with Principles 1 and 2. By incorporating the normalized discrepancy term $M^{\ast}(x,y_{w},y_{l})$, $\alpha$-DPO adaptively adjusts the margin between preferred and less preferred responses based on instance-specific differences, enhancing learning.
3 Theoretical Analysis of $\alpha$-DPO
In this section, we provide a theoretical analysis of $\alpha$-DPO by connecting its objective to the online SimPO loss. We also explore how $\alpha$-DPO manages the trade-off between alignment and diversity through KL divergence control.
3.1 Relation of $\alpha$-DPO to Online SimPO
In this section, we relate the $\alpha$-DPO objective to the online SimPO loss and demonstrate how our preference optimization method leads to a generalization bound for the policy model. We consider the following definitions:
Definition 3.1 (Online SimPO Loss).
Let $\pi_{\theta}$ represent the current policy and $\pi_{\theta_{\text{old}}}$ represent the policy from a previous iteration. The online SimPO loss is defined as:
$\mathcal{L}_{\text{SimPO}}^{\text{online}}(\pi_{\theta},\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\pi_{\theta_{\text{old}}}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$ (13)
where $\sigma(\cdot)$ is the sigmoid function, $\beta$ is a scaling parameter, $|y|$ represents the sequence length, and $\gamma$ is a margin parameter.
By applying importance sampling, this loss can be rewritten as:
$\mathcal{L}_{\text{SimPO}}^{\text{online}}(\pi_{\theta},\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\pi_{\text{ref}}}\left[w(y_{w},y_{l}|x)\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$ (14)
where the importance weight $w(y_{w},y_{l}|x)$ is given by:
$w(y_{w},y_{l}|x)=\frac{\pi_{\theta_{\text{old}}}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\cdot\frac{\pi_{\theta_{\text{old}}}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}.$ (15)
However, this importance weighting scheme can lead to undesirable results when $\pi_{\theta_{\text{old}}}(y_{l}|x)>\pi_{\text{ref}}(y_{l}|x)$, particularly for rejected responses. In such cases, the weight assigned to less preferred responses becomes disproportionately large, which is counterintuitive and can negatively impact the optimization process. To address this issue, we propose an alternative weighting scheme:
$w_{\text{corr}}(y_{w},y_{l}|x)=\frac{\pi_{\theta_{\text{old}}}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\cdot\frac{\pi_{\text{ref}}(y_{l}|x)}{\pi_{\theta_{\text{old}}}(y_{l}|x)}.$ (16)
This corrected weighting inversely scales the weight of overrepresented rejected responses, effectively reducing their influence in the expectation and improving the alignment between $\pi_{\theta}$ and $\pi_{\text{ref}}$. While this adjustment deviates from the standard importance sampling framework, it functions as a variance reduction technique that biases the estimator towards more desirable properties. Similar approaches have been explored in importance sampling literature to mitigate variance issues (Hanna et al., 2019; Patterson et al., 2021). We now establish that the $\alpha$-DPO objective serves as a lower bound on the online SimPO loss, justifying its use as a surrogate optimization objective.
Lemma 3.2 ($\alpha$-DPO provides a lower bound on the online SimPO loss).
For any policy model $\pi_{\theta}$ and reference model $\pi_{\text{ref}}$, there exists a sufficiently small $\alpha>0$ such that the following inequality holds:
$\mathcal{L}_{\text{$\alpha$-DPO}}(\pi_{\theta},\pi_{\text{ref}})\leq\mathcal{L}_{\text{SimPO}}^{\text{online}}(\pi_{\theta},\pi_{\text{ref}})+\mathbb{E}_{\pi_{\text{ref}}}\left[\alpha B\left(\log\sigma(A)-\sigma(A)+1\right)\right],$
where $A=\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma$, and $B=\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$.
Lemma 3.2 establishes that our objective $\mathcal{L}_{\text{$\alpha$-DPO}}(\pi_{\theta},\pi_{\text{ref}})$ provides a lower bound for the online SimPO loss $\mathcal{L}_{\text{SimPO}}^{\text{online}}(\pi_{\theta},\pi_{\text{ref}})$, up to a small approximation error dependent on $\alpha$. This implies that, under appropriate conditions (e.g., , as $\alpha\to 0$), optimizing $\mathcal{L}_{\text{$\alpha$-DPO}}$ can closely approximate the minimization of the online SimPO loss. Consequently, improvements in the $\alpha$-DPO objective lead to proportional enhancements in policy performance. Moreover, the lower-bounding property provides theoretical guarantees that the policy updated via $\alpha$-DPO will not perform worse than that optimized directly with the online SimPO loss, ensuring convergence to a well-generalized policy.
3.2 Balancing Alignment and Diversity via KL Divergence Control
@@
$\mathcal{L}_{\text{TDPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\delta(x,y_{w},y_{l})\right)\right],$ (17)
@@
$=\beta\sum_{t=1}^{|y_{l}|}\mathbb{E}_{z\sim\pi_{\text{ref}}}[\log\frac{\pi_{\text{ref}}(z|[x,y^{<t}_{l}])}{\pi_{\theta}(z|[x,y^{<t}_{l}])}]-\beta\sum_{t=1}^{|y_{w}|}\mathbb{E}_{z\sim\pi_{\text{ref}}}[\log\frac{\pi_{\text{ref}}(z|[x,y^{<t}_{w}])}{\pi_{\theta}(z|[x,y^{<t}_{w}])}]$ (19)
and $\mathbb{D}_{\text{SeqKL}}[x,y;\pi_{\text{ref}}||\pi_{\theta}]$ denotes the sequential KL divergence between $\pi_{\text{ref}}$ and $\pi_{\theta}$ along the sequence $y$ given $x$.
@@
Let $\delta(x,y_{w},y_{l})$ denote the difference in sequential KL divergences between the reference policy $\pi_{\text{ref}}$ and the policy $\pi_{\theta}$ along the sequences $y_{w}$ and $y_{l}$, respectively, defined as:
$\delta(x,y_{w},y_{l})=\beta\mathbb{D}_{\text{SeqKL}}[x,y_{l};\pi_{\text{ref}}||\pi_{\theta}]-\beta\mathbb{D}_{\text{SeqKL}}[x,y_{w};\pi_{\text{ref}}||\pi_{\theta}],$
If we approximate the sequential KL divergences using the log-probability ratios of the sequences, then $\delta(x,y_{w},y_{l})$ simplifies to:
$\delta(x,y_{w},y_{l})\approx\beta\left(\log\dfrac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\dfrac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right).$
Consequently, $\delta(x,y_{w},y_{l})$ is approximately equivalent to the margin term $M(x,y_{w},y_{l})$ in the $\alpha$-DPO objective.
The approximation leverages the assumption that the expectation over $z$ can be approximated by the log-probability ratios at the sequence level. Specifically, we recognize that the sequential KL divergence between $\pi_{\text{ref}}$ and $\pi_{\theta}$ along a sequence $y$ can be approximated by:
$\mathbb{D}_{\text{SeqKL}}[x,y;\pi_{\text{ref}}||\pi_{\theta}]$ $=\sum_{t=1}^{|y|}\mathbb{E}_{z\sim\pi_{\text{ref}}}\left[\log\dfrac{\pi_{\text{ref}}(z|x,y^{<t})}{\pi_{\theta}(z|x,y^{<t})}\right]\approx\log\dfrac{\pi_{\text{ref}}(y|x)}{\pi_{\theta}(y|x)}.$
This approximation reduces computational complexity by operating at the sequence level rather than the token level, making it particularly advantageous when dealing with long sequences or when the reference policy $\pi_{\text{ref}}$ is not well-calibrated at the token level. Applying this approximation to both $y_{w}$ and $y_{l}$, the difference $\delta(x,y_{w},y_{l})$ simplifies to the difference of log-probability ratios, thereby establishing the equivalence with the margin term in $\alpha$-DPO.
Lemma 3.3 highlights that the margin term $\delta(x,y_{w},y_{l})$, which represents the sequential KL divergence difference between preferred and rejected responses, can be directly mapped to the term $M(x,y_{w},y_{l})$ in $\alpha$-DPO. This mapping underscores the theoretical connection between the two approaches in terms of alignment control. While TDPO operates at the token level and provides fine-grained control, $\alpha$-DPO offers greater computational efficiency by operating at the sequence level without sacrificing performance. Moreover, the sequence-level approximation enhances robustness to token-level noise in $\pi_{\text{ref}}$, making $\alpha$-DPO particularly suited for scenarios where the reference policy may not be perfectly aligned. Refer to Appendix B.4, where we compare TDPO and $\alpha$-DPO.

### v4

# AlphaDPO: Adaptive Reward Margin for Direct Preference Optimization

@@ Abstract
Aligning large language models (LLMs) with human preferences requires balancing policy optimization with computational stability. While recent offline methods like DPO and SimPO bypass reinforcement learning’s complexity, they face critical limitations: DPO relies on static reference models that degrade with policy updates, and SimPO assumes a uniform target reward margin that ignores instance-wise preference strength. We propose AlphaDPO, an adaptive preference optimization framework that dynamically reparameterizes the reference distribution to address these issues. Our key innovation lies in an implicit reference model $\hat{\pi}_{\text{ref}}\propto U(y|x)(\pi_{\theta}/\pi_{\text{ref}})^{\alpha}$, which interpolates between policy-driven specialization and uniform exploration while enabling instance-adaptive reward margins. Theoretically, we prove AlphaDPO implicitly controls sequential KL divergence between iterative policy updates, ensuring stability even with poorly calibrated reference models. Empirically, AlphaDPO achieves state-of-the-art performance on AlpacaEval 2 (58.7% LC win rate) and Arena-Hard (35.7% win rate) across Mistral2-7B, Llama3-8B, and Gemma2-9B, demonstrating robust alignment without multi-stage training. Our work establishes adaptive reference reparameterization as a principled mechanism for preference optimization. The code is available at https://github.com/junkangwu/alpha-DPO.
Machine Learning, ICML
@@ 1 Preliminaries
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (4)
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right].$
@@
${\mathcal{L}}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (5)
$\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$
@@ 2 Method
In this section, we establish a unified framework that connects DPO and SimPO (Section 2.1), highlighting the critical role of the reference model in preference optimization. We then introduce AlphaDPO (Section 2.2), a new preference optimization algorithm that synergizes the strengths of both DPO and SimPO.
@@ 2.1 A Common Framework for DPO and SimPO
$\mathcal{L}(\pi_{\theta};U)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (6)
$\left[\log\sigma\left(\beta\left(\log\pi_{\theta}(y_{w}|x)-\log\pi_{\theta}(y_{l}|x)\right)-\gamma\right)\right],$
@@
$\mathcal{L}_{\text{LN}}(\pi_{\theta};U)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (7)
$\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right].$
@@
Remark 2.2.
Why do winning and losing responses have different probabilities under a uniform policy? Although both winning and losing responses are sampled from the same policy model, their selection probabilities diverge through the reward model’s scoring mechanism. Specifically, winning response ($y_{w}$) are those assigned higher scores, while losing response ($y_{l}$) receive lower scores. Even under a uniform policy, these selections remain distinct because their selection probabilities are tied to their relative scores, not to the underlying uniform distribution over the response space.
Theorem 2.1 establishes a unified framework for DPO and SimPO by showing that replacing the reference model $\pi_{\text{ref}}$ in DPO with a uniform distribution $U$ reduces the DPO loss to the SimPO loss, up to a constant term $\gamma$. This reveals that SimPO is essentially DPO with a uniform reference model. Consequently, the term $\beta\left(\log\pi_{\text{ref}}(y_{w}|x)-\log\pi_{\text{ref}}(y_{l}|x)\right)$ collapses to a constant, reflecting the difference in selection probabilities induced by the scoring mechanism. This emphasizes the pivotal role of the implicit reference model in preference optimization.
r  Refer to caption
Figure 1: $\log\pi_{\mathrm{ref}}(y_{w}|x)-\log\pi_{\mathrm{ref}}(y_{l}|x)$ along the training steps. The random fluctuations suggest $\pi_{\mathrm{ref}}$ struggles to distinguish between the preferred and less preferred responses.
@@
Limitations of SimPO: While SimPO simplifies the loss function by using a constant offset $\gamma$, this one-size-fits-all approach overlooks the variability inherent in different data instances. The fixed $\gamma$ across all training samples could lead to suboptimal performance, especially in the presence of noise or inconsistencies in the data. Moreover, completely discarding reference model $\pi_{\text{ref}}$ may eliminate potentially useful prior knowledge about language structure and semantic relationships that could help discriminate between similar response pairs.
2.2 Proposed Method: AlphaDPO
@@
Deriving the AlphaDPO Objective. Starting from the standard Reinforcement Learning (RL) objective for preference optimization, we redefine the reference model $\pi_{\text{ref}}$ as an implicit reference model $\hat{\pi}_{\text{ref}}$, formulated as:
$\hat{\pi}_{\text{ref}}(y|x)\propto U(y|x)\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\alpha},\vspace{-15pt}$ (8)
where $\alpha$ is a hyperparameter that determines the extent to which the policy model $\pi_{\theta}$ modifies the reference model $\pi_{\text{ref}}$, and $U(y|x)$ represents a uniform distribution providing a stable baseline. This formulation shares structural similarities with the weak-to-strong preference optimization framework (Zhu et al., 2025), where the aligned strong model $\pi_{r}^{\text{strong}}$ is expressed as:
$\vspace{-15pt}\pi_{r}^{\text{strong}}(y\mid x)\propto\pi_{\text{ref}}^{\text{strong}}(y\mid x)\left(\frac{\pi_{r}^{\text{weak}}(y\mid x)}{\pi_{\text{ref}}^{\text{weak}}(y\mid x)}\right)^{\alpha}.$ (9)
Both formulations employ model ratios to dynamically adjust the reference distribution — $\frac{\pi_{\theta}}{\pi_{\text{ref}}}$ in Equation 8 and $\frac{\pi_{r}^{\text{weak}}}{\pi_{\text{ref}}^{\text{weak}}}$ in Equation 9. Specifically, $\hat{\pi}_{\text{ref}}$ interpolates between existing methods, controlled by the smoothness parameter $\alpha$: i) When $\alpha=0$, $\hat{\pi}_{\text{ref}}$ reduces to the uniform distribution $U(y|x)$, equivalent to the assumption in SimPO. ii) When $\alpha=1$, $\hat{\pi}_{\text{ref}}$ fully incorporates the proportionality $\frac{\pi_{\theta}}{\pi_{\text{ref}}}$, aligning with DPO.
The motivation behind this redefinition is to overcome the potential suboptimality of both SimPO and DPO by learning instance-specific adjustments while preserving useful prior knowledge. Rather than relying on either a fixed reference model or a constant margin, $\hat{\pi}_{\text{ref}}$ creates a flexible framework that can better capture varying degrees of preference strength across different instances. Further details are provided in Appendix C.
Substituting $\hat{\pi}_{\text{ref}}$ into the original DPO loss function, we obtain the AlphaDPO objective:
$\mathcal{L}_{\text{AlphaDPO}}(\pi_{\theta},\pi_{\text{ref}})$ (10)
$=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}-\beta\log\frac{\hat{\pi}_{\text{ref}}(y_{w}|x)}{\hat{\pi}_{\text{ref}}(y_{l}|x)}\right)\right]$
$=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right)-\left[\gamma+\alpha M(x,y_{w},y_{l})\right]\right)\right],$
@@
Length-normalized Reward Formulation: Inspired by SimPO, we incorporate length normalization into our method. This adjustment ensures that rewards are scaled appropriately with respect to the length of the sequences, thereby stabilizing the training process. As demonstrated in our experiments (cf. Appendix D.2), we also confirmed that even without length normalization, our method remains effective and continues to show performance improvements.
@@
Final Objective: Incorporating the above considerations, the final AlphaDPO loss function becomes:
$\mathcal{L}_{\text{AlphaDPO}}(\pi_{\theta},\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (13)
$\left[\log\sigma\left(u(x,y_{w},y_{l})-\operatorname{sg}\left[\gamma+\alpha M^{\ast}(x,y_{w},y_{l})\right]\right)\right],$
@@
where $u(x,y_{w},y_{l})=\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log{\pi_{\theta}(y_{l}|x)}$. This formulation ensures balanced influence between the policy and reference models, aligning with Principles 1 and 2. By incorporating the normalized discrepancy term $M^{\ast}(x,y_{w},y_{l})$, AlphaDPO adaptively adjusts the margin between preferred and less preferred responses based on instance-specific differences, enhancing learning.
3 Theoretical Analysis of AlphaDPO
@@
$\mathcal{L}_{\text{TDPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (14)
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\delta(x,y_{w},y_{l})\right)\right],$
@@
$=\beta\sum_{t=1}^{|y_{l}|}\mathbb{E}_{z}[\log\frac{\pi_{\text{ref}}(z|[x,y^{<t}_{l}])}{\pi_{\theta}(z|[x,y^{<t}_{l}])}]-\beta\sum_{t=1}^{|y_{w}|}\mathbb{E}_{z}[\log\frac{\pi_{\text{ref}}(z|[x,y^{<t}_{w}])}{\pi_{\theta}(z|[x,y^{<t}_{w}])}].$ (16)
Here, $z\sim\pi_{\text{ref}}$ and $\mathbb{D}_{\text{SeqKL}}[x,y;\pi_{\text{ref}}||\pi_{\theta}]$ denotes the sequential KL divergence between $\pi_{\text{ref}}$ and $\pi_{\theta}$ along the sequence $y$ given $x$.
Below we present a lemma establishing theoretical connections between TDPO and AlphaDPO w.r.t. the margin terms.
@@
Consider the margin term $\delta(x,y_{w},y_{l})$ that represents the difference in sequential KL divergences between a reference policy $\pi_{ref}$ and policy $\pi_{\theta}$ for preferred sequence $y_{w}$ and rejected sequence $y_{l}$, conditioned on input x:
$\delta(x,y_{w},y_{l})$ $=\beta\mathbb{D}_{\text{SeqKL}}[x,y_{l};\pi_{\text{ref}}||\pi_{\theta}]$ (17)
$\quad-\beta\mathbb{D}_{\text{SeqKL}}[x,y_{w};\pi_{\text{ref}}||\pi_{\theta}],$
Under the assumption that the sequential KL divergence can be approximated by the log-likelihood ratio of complete sequences, the margin term $\delta(x,y_{w},y_{l})$ admits the following approximation:
$\delta(x,y_{w},y_{l})$ $\approx\beta\left(\log\dfrac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\dfrac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
$=M(x,y_{w},y_{l}),$
where $M(x,y_{w},y_{l})$ is the margin term in the AlphaDPO objective.
The proof follows from the sequence-level approximation of the KL divergence between $\pi_{\text{ref}}$ and $\pi_{\theta}$ along a sequence $y$, where:
$\mathbb{D}_{\text{SeqKL}}[x,y;\pi_{\text{ref}}||\pi_{\theta}]$ $=\sum_{t=1}^{|y|}\mathbb{E}_{z\sim\pi_{\text{ref}}}\left[\log\dfrac{\pi_{\text{ref}}(z|x,y^{<t})}{\pi_{\theta}(z|x,y^{<t})}\right]$
$\approx\log\dfrac{\pi_{\text{ref}}(y|x)}{\pi_{\theta}(y|x)}.$
Applying this approximation to both $y_{w}$ and $y_{l}$, the difference $\delta(x,y_{w},y_{l})$ simplifies to the difference of log-probability ratios, thereby establishing the equivalence with the margin term in AlphaDPO.
Lemma 3.1 highlights that the margin term $\delta(x,y_{w},y_{l})$, which represents the sequential KL divergence difference between preferred and rejected responses, can be directly mapped to the term $M(x,y_{w},y_{l})$ in AlphaDPO. This mapping underscores the theoretical connection between the two approaches in terms of alignment control. While TDPO operates at the token level and provides fine-grained control, AlphaDPO offers greater computational efficiency by operating at the sequence level without sacrificing performance (cf. Appendix Table 6). Moreover, the sequence-level approximation enhances robustness to token-level noise in $\pi_{\text{ref}}$, making AlphaDPO particularly suited for scenarios where the reference policy may not be perfectly aligned.
Advantages of the margin term $\delta(x,y_{w},y_{l})$. The core contribution of TDPO lies in introducing the margin term $r(x,y_{w})-r(x,y_{l})-\delta(x,y_{w},y_{l})$, which is similar to DPO with an offset (Amini et al., 2024) and helps control the KL divergence. In contrast, AlphaDPO generalizes this approach by replacing $\delta(x,y_{w},y_{l})$ with $M(x,y_{w},y_{l})$, a more flexible margin term inspired by SimPO. Appendix Table 6 supports this with performance comparisons between TDPO and AlphaDPO. These findings illustrate: i) Adding an offset to DPO and its variants is a robust strategy, applicable to both TDPO and AlphaDPO. ii) The choice of offset is critical — the sequence-level margin term $M(x,y_{w},y_{l})$ is particularly effective when applied to potentially unreliable reference models.